A lift is reported to operate at 75% efficiency. If the lift can raise a load through a height of 30m using a 230V, 15A supply, calculate how long it will take to raise a load of 1000Kg.


1
Expert's answer
2021-11-09T10:58:41-0500

"g = 9.8\frac{m}{s^2}"

"h = 30m"

"m =1000kg"

"\eta=75 \%"

"I = 15A"

"U= 230V"

"\Delta EP = mgh = 9.8*30*1000= 294000\ J"

"W=\Delta EP= 294000\ J"

"\eta =\frac{P_{out}}{P_{inp}}*100\%"

"P_{inp}= IU= 230*15= 3450\ W"

"P_{out}=\frac{P_{inp}\eta}{100\%}=2587.5\ W"

"W =P_{out}t"

"t =\frac{W}{P_{out}}=\frac{294000}{2587.5}=113.62\ s"


"\text{Answer: }t = 113.62\ s"